# Information Discovery in e-Commerce

### v1

@@ 8.1 Conclusion
From the large number of user engagements on e-commerce platforms a large amount of information can be inferred. Information discovery has increasingly been attracting attention from information retrieval community. The aim of this survey has been to give a broad overview of information discovery in e-commerce portals. Our overview has included methods about user behavior modeling, search, recommendation, question answering, and dialogue systems in e-commerce.
@@
Our strategy with this survey has been to provide a broad coverage of research directions about information discovery in e-commerce. Although we have tried our best to provide all key approaches in each direction as much as possible, the amount of technical details is limited. For areas that are broad enough to have their own survey, we only focus on key publications and provide structure and pipelines for each direction. Additionally, we only focus on areas that are relevant to information retrieval research; studies in other areas relevant to e-commerce, such as supply chains and computational advertising, are ignored in this survey.
@@
We have briefly presented information discovery research in e-commerce scenarios. In the introduction, we summarized the outline and topics covered in this survey, followed by a chapter with basic concepts and key definitions in Chapter 2. Then we introduced preliminaries about e-commerce interfaces and users in Chapter 3. We formulated concepts of e-commerce infrastructures and summarized studies about information seeking via e-commerce interfaces in Section 3.1. We investigated e-commerce information components, i.e., titles, product descriptions, and reviews, and detailed characteristics of consumer behaviors in e-commerce portals. We introduced studies into e-commerce user analyses concerning multiple behaviors, including clicks, purchases, engagements, and post-clicks, on e-commerce platforms.
@@
The core of this survey is organized around five directions: e-commerce user modeling, e-commerce search, e-commerce recommendation, e-commerce QA, and e-commerce dialogue systems. We have detailed each of these in four chapters (i.e., Chapter 456, and 7). Each chapter starts with an overview of the main direction discussed in the chapter, with characteristics and subtasks. After that, key research studies of each subtasks were demonstrated with some level of detail. We discussed emerging research directions at the end of each key component.
@@
In particular, in Chapter 4, we introduced approaches to user modeling and profiling for e-commerce applications. We divided this chapter into two main components: user behavior modeling and user profiling. We provided a summary of studies on modeling these e-commerce user behavior, analyzed research on user profiling in e-commerce, and discussed emerging directions on user modeling in e-commerce.
@@
In Chapter 5, we focused on search technologies in e-commerce platforms. We provided the characteristics of e-commerce search and divided research studies based on matching strategies and ranking technologies for e-commerce search scenarios, respectively. We presented approaches aiming for studying matching strategies for e-commerce search. And we studied research approaches on ranking technologies for e-commerce search. Emerging research directions were discussed at the end of the chapter.
@@
In Chapter 6, we introduced the most prominent approaches on e-commerce recommendation methods. We summarized the key characteristics of e-commerce recommendation, towards which a two-stage framework was developed that contains candidate retrieval and candidate ranking, forming the mainstream solution for e-commerce recommender systems. We reviewed models developed for the two stages and detailed mainstream learning methods for optimizing model parameters to provide a complete view of e-commerce recommender systems.
@@
Chapter 7 detailed research methods for question answering and dialogue systems in e-commerce. We merged question answering and dialogue systems into a single chapter as most research background and approaches are shared in these two directions. We reviewed previous work on question answering and then demonstrated the characteristics of e-commerce QA. For e-commerce question answering (QA), we described studies both on extractive QA and generative QA.. For e-commerce dialogue systems, we demonstrated the patterns of e-commerce dialogue systems, especially about task-oriented dialogue systems, knowledge-grounded conversations, and empathetic dialogue systems. We discussed emerging research directions around QA and dialogue agents in Section 7.3.
@@ 8.2 Outlook
Information discovery is increasingly a mixed initiative scenario, where users and e-commerce platforms take turns. As described in the previous chapters, the research presented on information discovery in e-commerce has been addressed from six angles: infrastructures, user modeling, search, recommendation, QA, and dialogue systems. As we summarized at the end of each chapter, a broad variety of emerging research has also been motivated following each angle. For user modeling, we consider three research topics as emerging directions on user modeling and profiling: graph learning for user behavior modeling, dynamic user behavior modeling and profiling, and multi-modal user profiling. For emerging directions on e-commerce search, we focus on applications for multi-modal e-commerce search and ranking. Online learning to rank technologies also provide insights. We also foresee the development of new learning theories that will improve e-commerce search and ranking performance in future. We divide emerging directions on e-commerce recommendation into three directions: (i) reasoning recommendation and explanations; (ii) conversational recommendation; and (iii) unifying recommendation and search. In our view, future work on e-commerce language processing should includes generating explainable reasons for search and recommendation, improving robustness e-commerce question answering, and improving conversational e-commerce search and recommendation.
@@ 8.2.1 Four directions
Conversational search refers to a novel search paradigm using interactions between users and search engines. As we have discussed in Chapter 5, conversational search is increasingly receiving more attention in the IR community. Different from the traditional query-aware search paradigm, conversational search allows users to express their information need by directly conducting conversations with search engines. Recent studies start to apply conversational search to online shopping scenarios as it is able to provide natural, adaptive and interactive shopping experience for consumers [xiao2021end]. In e-commerce, conversational search faces two challenges: imperfect product attribute schemas and product knowledge. The former exists as product attributes link lengthy multi-turn utterances with products in conversational search systems, whereas the derives from the lack of manually labelling in benchmark datasets. Core tasks in conversational search, e.g., search intent detection, action prediction, query selection, passage selection, and response generation [ren2021wizard], also provide insights in future work about e-commerce conversational search systems.
@@
Conversational recommendation refers to a recommendation system that can elicit the dynamic preferences of users and take actions based on their current needs through real-time multi-turn interactions. As we have discussed in Chapter 6, conversational recommendation is becoming an emerging research direction for e-commerce recommendation. Integrating more accurate domain-specific knowledge to promote the recommendation and conversation is still a challenging problem in conversational recommendation [chen2020towards]. Moreover, as with conversational search, current studies on conversational recommendation suffer from the lack of manually labelled data in benchmark datasets. Recent studies on empathetic dialogue systems reveal that there exist some kind of dependency between commonsense knowledge and emotional preference [li2020empathetic, siro-2022-understanding]. Hence, a conversational recommender system jointly combining emotion detection and knowledge exploration is worth studying in future.
@@
Multi-modal information can be widely observed in many e-commerce scenarios, e.g., user behaviors, search, recommendation, and dialogue systems. Most previous e-commerce information discovery approaches are constructed only based on text understanding and retrieval; addressing multi-modal information, e.g., images and videos, still appears to be difficult. In future work, it is more and more important to tackle challenges about multi-modality in e-commerce scenarios. Multimedia technologies focusing on integrating various types of modalities are expected to help to understand those multi-modal information for various types of e-commerce applications. Moreover, it is interesting to explore multi-modal generation through more powerful generative deep neural networks in e-commerce review generation, question answering, and dialogue systems.
@@
Large scale generative models have the potential to significantly enhance various e-commerce information discovery applications, such as search, recommendation, and conversational AI. Transformer-based pre-trained language models like BERT have already proven effective in both search and recommendation tasks in e-commerce. More recently, large language models (LLMs) based on auto-regressive mechanisms, such as T5 and GPT, have demonstrated promising capabilities in understanding and generating human-like information, making them valuable for e-commerce contexts. Moreover, while traditional two-stage paradigms (i.e., retrieval followed by re-ranking) have been widely used in e-commerce search and recommendation scenarios. They face two limitations: (i) heterogeneous modules with different optimization objectives may lead to sub-optimal performance; and (ii) a large document index is needed which may come with substantial memory and computational requirements. This has motivated research into end-to-end solutions using generative models. Recent studies on generative retrieval, such as DSI [tay2022transformer], have shown encouraging performance on several information retrieval benchmarks, suggesting that exploring generative models for end-to-end e-commerce search and recommendation could be a promising direction for future research.
@@ 8.2.2 Beyond accuracy
Besides the three directions for future work listed above, we also consider the following important issues when it comes to information discovery tasks in e-commerce: fairness, trustworthiness, and explainability [derijke-2023-beyond].
@@
Recently, the problem of bias has attracted considerable attention in the IR community, in multiple contexts, e.g., for user behavior modeling, profiling, ranking, and recommendation. To address the bias problem, fairness is considered as a significant metric during the optimization procedure. Early on, fairness was studied from the perspective of information exposure regarding sensitive attributes such as gender and race [singh2018fairness]. Fairness in IR also focuses on how to let different items receive equal exposure, or exposure proportional to their utility or impact, depending on which exposure distribution is considered to be fair by the system [morik2020controlling, chen2020bias]. In e-commerce, ranking-based interfaces are quite common in various scenarios, which makes the fairness issue a growing influence role for information discovery. Future work on interactive fairness-aware reranking can be helpful for debiasing user modeling, search, recommendation, and answer generation in e-commerce platforms. Also, knowledge-based and dynamic fairness-aware methods are able to address more real-world challenges. The trade-off between accuracy and fairness is of importance in e-commerce search and recommendation scenarios, where equally treating different groups has been shown to sacrifice the performance. To address this problem, an important research direction is to understand the dimensions of causality and design fairness-aware algorithms.
@@
Fake news and fake information are is increasingly widespread. It is now viewed as one of the greatest threats on the web [zhou2020survey]. As we have discussed in Chapter 7, spam and fake reviews and answers are widely observed in many e-commerce platforms. Therefore, pursuing trustworthiness has become an important issue in e-commerce question answering and dialogue systems. Recent studies aim to distinguish spam or fake reviews in online review systems via graph neural networks [kaghazgaran2018combating, rao2020xfraud, liu2020alleviating, dou2020enhancing]. Textual generation methods based on deep neural networks have been applied to fake reviews or spam generation by camouflaged fraudsters. Thus, distinguishing the authenticity of e-commerce information is a challenging task. Other patterns in the reviews, e.g., sentiments and emotions, can be applied to improve the detection. Also, investigating inconsistency problems under multiple domains provides new avenues of research.
@@
Explainability in e-commerce aims to answer the question about why we receive a specific ranking, recommendation, or answer results. The task of explainability can be divided into to types: the explainability of the learning models and the explainability of the results. The former aims to provide more transparent learning details for the proposed methods, whereas the latter focuses on provide more explainable results to various application scenarios, e.g., search, recommendation, and question answering, etc. Explainability methods have been shown to be effective for enhancing the e-commerce search and recommendation [zhang2018towards, liu2020keywords]. For future work, it is important to evaluate if and how users are satisfied with the explanations generated from an e-commerce system, especially as these are increasingly conversational in nature. Generating coherent and naturally-sounded explanations based on a sequence of reasoning steps (including search or recommendation system output) is still difficult.
@@
In a broad sense, researchers in the broader communities, e.g., AI and operational research, have also realized the importance of information discovery in e-commerce, which aims to address a wide range of information discovery problems in e-commerce applications. This survey would be useful for both academic and industrial researchers who are working on either information retrieval or e-commerce. As we have discussed in our survey, the number of emerging research studies is increasing in recent years. We believe that this is only the beginning of the whole research task. The recent launch of a dedicated product search track at TREC seems to confirm this. [^1] https://trec-product-search.github.io The quantity of the work described in this survey and the steady pace of publications in the field together with the open research follow-ups indicate a promising future ahead. There is a lot of future work to be done.
@@ Acknowledgements.
This survey grew out of a tutorial “Information Discovery in e-Commerce” taught at SIGIR 2018. We thank the audience for their feedback and questions. We also thank our colleagues Mozhdeh Ariannezhad, Hongshen Chen, Jiawei Chen, Zhumin Chen, Zhuoye Ding, Yue Feng, Yulong Gu, Shuyu Guo, Ziyi Guo, Maria Heuss, Mariya Hendriksen, Na Huang, Sami Jullien, Barrie Kersbergen, Jiahuan Lei, Dongdong Li, Ming Li, Xiang Li, Xinyi Li, Zhenyang Li, Xiaozhong Liu, Hengliang Luo, Si Luo, Yougang Lyu, Jun Ma, Yao Ma, Pengjie Ren, Fatemeh Sarvi, Sebastian Schelter, Xinlei Shi, Clemencia Siro, Hongye Song, Olivier Sprangers, Changlong Sun, Fei Sun, Weiwei Sun, Jiliang Tang, Jingang Wang, Shuaiqiang Wang, Xuepeng Wang, Zihan Wang, Long Xia, Xin Xin, Zhen Zhang, Jiashu Zhao, Xiangyu Zhao, and Yihong Zhao for help, feedback, and inspiration. We thank our editors Yiqun Liu and Mark Sanderson for support and valuable feedback. This research was supported by Baidu.com, JD.com, Alibaba DAMO Academy, Meituan, and by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v3

@@ 8.1 Conclusion
From the large number of user engagements on e-commerce platforms a large amount of information can be inferred. The aim of this survey has been to give a broad overview of information discovery in e-commerce portals. Our overview has included methods about user behavior modeling, search, recommendation, question answering, and dialogue systems in e-commerce.
@@
Our strategy with this survey has been to provide a broad coverage of research directions about information discovery in e-commerce. Although we have tried our best to provide all key approaches in each direction as much as possible, the amount of technical details is limited. For areas that are broad enough to have their own survey, we have only focused on key publications and provide structure and pipelines for each direction. Additionally, we only focus on areas that are relevant to information retrieval research; studies in other areas relevant to e-commerce, such as supply chains and computational advertising, are ignored in this survey.
@@
In our introduction, we summarized the outline and topics covered in this survey, followed by a description of basic concepts and key definitions in Section 2. Then we introduced preliminaries about e-commerce interfaces and users in Section 3. We formulated concepts of e-commerce infrastructures and summarized studies about information seeking via e-commerce interfaces. We investigated e-commerce information components, i.e., titles, product descriptions, and reviews, and detailed characteristics of consumer behaviors in e-commerce portals. We introduced studies into e-commerce user analyses concerning multiple behaviors, including clicks, purchases, engagements, and post-clicks, on e-commerce platforms.
@@
The core of this survey is organized around five directions: e-commerce user modeling, e-commerce search, e-commerce recommendation, e-commerce QA, and e-commerce dialogue systems. We have detailed each of these in four sections (i.e., Section 456, and 7). Each section starts with an overview of the main direction discussed in the section, with characteristics and subtasks. After that, key research studies of each subtasks were demonstrated with some level of detail. We discussed emerging research directions at the end of each key component.
@@
In particular, in Section 4, we introduced approaches to user modeling and profiling for e-commerce applications. We divided this section into two main components: user behavior modeling and user profiling. We provided a summary of studies on modeling these e-commerce user behavior, analyzed research on user profiling in e-commerce, and discussed emerging directions on user modeling in e-commerce.
@@
In Section 5, we focused on search technologies in e-commerce platforms. We provided the characteristics of e-commerce search and divided research studies based on matching strategies and ranking technologies for e-commerce search scenarios, respectively. We presented approaches aiming for studying matching strategies for e-commerce search. And we studied research approaches on ranking technologies for e-commerce search. Emerging research directions were discussed at the end of the section.
@@
In Section 6, we introduced the most prominent approaches to e-commerce recommendation methods. We summarized the key characteristics of e-commerce recommendation, towards which a two-stage framework was developed that contains candidate retrieval and candidate ranking, forming the mainstream solution for e-commerce recommender systems. We reviewed models developed for the two stages and detailed mainstream learning methods for optimizing model parameters to provide a complete view of e-commerce recommender systems.
@@
Section 7 detailed research methods for question answering and dialogue systems in e-commerce. We addressed question answering and dialogue systems in a single section as most research background and approaches are shared between these two directions. We reviewed previous work on question answering and then demonstrated the characteristics of e-commerce QA. For e-commerce question answering (QA), we described studies both on extractive QA and generative QA. For e-commerce dialogue systems, we demonstrated the patterns of e-commerce dialogue systems, especially about task-oriented dialogue systems, knowledge-grounded conversations, and empathetic dialogue systems. We discussed emerging research directions around QA and dialogue agents in Section 7.3.
@@ 8.2 Outlook
Information discovery is increasingly a mixed initiative scenario, where users and e-commerce platforms take turns. As described in the previous sections, the research presented on information discovery in e-commerce has been addressed from six angles: infrastructures, user modeling, search, recommendation, QA, and dialogue systems. As we summarized at the end of each section, a broad variety of emerging research has also been motivated following each angle. For user modeling, we consider three research topics as key emerging directions: graph learning for user behavior modeling, dynamic user behavior modeling and profiling, and multi-modal user profiling. For emerging directions in e-commerce search, we focus on applications for multi-modal e-commerce search and ranking. Online learning to rank technologies also provide key insights. We also foresee the development of new learning theories that will improve e-commerce search and ranking performance in the future. We divide emerging directions on e-commerce recommendation into three directions: (i) reasoning, recommendation, and explanations; (ii) conversational recommendation; and (iii) unifying recommendation and search. In our view, future work on e-commerce language processing should include generating explainable reasons for search and recommendation, improving the robustness of e-commerce question answering, and improving conversational e-commerce search and recommendation.
@@ 8.2.1 Four directions
Conversational search refers to a novel search paradigm using multiple interactions between users and search engines. As we have discussed in Section 5, conversational search is increasingly receiving more attention in the IR community. Different from the traditional query-aware search paradigm, conversational search allows users to express their information need by directly conducting conversations with search engines. More recent studies have begun to apply conversational search to online shopping scenarios as it is able to provide a natural, adaptive and interactive shopping experience for consumers [xiao2021end]. In e-commerce, conversational search faces two challenges: imperfect product attribute schemas and product knowledge. The former exists as product attributes link lengthy multi-turn utterances with products in conversational search systems, whereas the latter derives from the lack of manually labelling in benchmark datasets. Core tasks in conversational search, e.g., search intent detection, action prediction, query selection, passage selection, and response generation [ren2021wizard], also provide insights in future work about e-commerce conversational search systems.
@@
Conversational recommendation refers to recommendation systems that can elicit the dynamic preferences of users and take actions based on their current needs through real-time multi-turn interactions. As we have discussed in Section 6, conversational recommendation is an emerging direction in e-commerce recommendation. Integrating more accurate domain-specific knowledge to promote the recommendation and conversation is a challenging problem in conversational recommendation [chen2020towards]. Moreover, as with conversational search, current studies on conversational recommendation suffer from a lack of manually labelled data in benchmark datasets. Recent studies on empathetic dialogue systems reveal that there exist some kind of dependency between commonsense knowledge and emotional preference [li2020empathetic, siro-2022-understanding]. Hence, conversational recommender systems that jointly combine emotion detection and knowledge exploration are worth studying in future.
@@
Multi-modal information can be widely observed in many e-commerce scenarios, e.g., user behavior, search, recommendation, and dialogue systems. Most previous e-commerce information discovery approaches are constructed only based on text understanding and retrieval; addressing multi-modal information, e.g., images and videos, still appears to be difficult. In future work, it is more and more important to tackle challenges about multi-modality in e-commerce scenarios. Multimedia technologies focusing on integrating various types of modalities are expected to help to understand those multi-modal information for various types of e-commerce applications. It is interesting to explore multi-modal generation through powerful generative deep neural networks in e-commerce review generation, question answering, and dialogue systems.
@@
Large-scale generative models have the potential to significantly enhance various e-commerce information discovery applications, such as search, recommendation, and conversational AI. Transformer-based pre-trained language models like BERT have already proven effective in both search and recommendation tasks in e-commerce. More recently, large language models (LLMs) based on auto-regressive mechanisms, such as T5 and GPT, have demonstrated promising capabilities in understanding and generating human-like information, making them valuable for e-commerce contexts. Moreover, while traditional two-stage paradigms (i.e., retrieval followed by re-ranking) have been widely used in e-commerce search and recommendation scenarios. They face two limitations: (i) heterogeneous modules with different optimization objectives may lead to sub-optimal performance; and (ii) a large document index is needed which may come with substantial memory and computational requirements. This has motivated research into end-to-end solutions using generative models. Recent studies on generative retrieval, such as DSI [tay2022transformer], have shown encouraging performance on several information retrieval benchmarks, suggesting that exploring generative models for end-to-end e-commerce search and recommendation could be a promising direction for future research.
@@ 8.2.2 Beyond accuracy
Besides the directions for future work listed above, we also consider the following important issues when it comes to information discovery tasks in e-commerce: fairness, trustworthiness, and explainability [roegiest-2019-facts-ir, derijke-2023-beyond].
@@
Recently, the problem of bias has attracted considerable attention in the IR community, in multiple contexts, e.g., for user behavior modeling, profiling, ranking, and recommendation. To address the bias problem, fairness is considered as a significant metric during the optimization procedure. Early on, fairness was studied from the perspective of information exposure regarding sensitive attributes such as gender and race [singh2018fairness]. Fairness in IR also focuses on how to let different items receive equal exposure, or exposure proportional to their utility or impact, depending on which exposure distribution is considered to be fair by the system [morik2020controlling, chen2020bias]. In e-commerce, ranking-based interfaces are quite common in various scenarios; hence, fairness is a matter of great importance to information discovery in e-commerce. Future work on interactive fairness-aware reranking can be helpful for debiasing user modeling, search, recommendation, and answer generation in e-commerce platforms [sarvi-2022-understanding]. Also, knowledge-based and dynamic fairness-aware methods are able to address more real-world challenges. The trade-off between accuracy and fairness is of importance in e-commerce search and recommendation scenarios, where equally treating different groups has been shown to sacrifice the performance [ariannezhad-2023-personalized]. To address this problem, an important research direction is to understand the dimensions of causality and design fairness-aware algorithms.
@@
Fake news and fake information are is increasingly widespread. It is now viewed as one of the greatest threats on the web [zhou2020survey]. As we have discussed in Section 7, spam and fake reviews and answers are widely observed in many e-commerce platforms. Therefore, pursuing trustworthiness has become an important issue in e-commerce question answering and dialogue systems. Several studies distinguish spam or fake reviews in online review systems via graph neural networks [kaghazgaran2018combating, rao2020xfraud, liu2020alleviating, dou2020enhancing]. Textual generation methods based on deep neural networks have been applied to fake reviews or spam generation by camouflaged fraudsters. Thus, distinguishing the authenticity of e-commerce information is a challenging task. Other patterns in the reviews, e.g., sentiments and emotions, can be applied to improve the detection. Also, investigating inconsistency problems under multiple domains provides new avenues of research.
@@
Explainability in e-commerce aims to answer the question about why we receive a specific ranking, recommendation, or answer result. The task of explainability can be divided into explainability of the learning models and explainability of the results. The former aims to provide more transparent learning details for the proposed methods, whereas the latter focuses on provide more explainable results to various application scenarios, e.g., search, recommendation, and question answering, etc. Explainability methods have been shown to be effective for enhancing the e-commerce search and recommendation [zhang2018towards, liu2020keywords]. For future work, it is important to evaluate if and how users and other stakeholders are satisfied with the explanations generated from an e-commerce system, especially as these are increasingly conversational in nature [lucic-2021-multistakeholder]. Generating coherent, faithful, and naturally-sounding explanations based on a sequence of reasoning steps (including search or recommendation system output) is still difficult.
@@
As we have shown in this survey, the number of research studies in the area of information discovery for e-commerce is increasing rapidly. We believe that this is only the beginning. The recent launch of a dedicated product search track at TREC seems to confirm this. [^1] https://trec-product-search.github.io The volume of the work described in this survey and the steady pace of publications in the field, together with the arrival of open research challenges indicate a promising future ahead. A lot remains to be done.
@@ Acknowledgements.
This survey grew out of a tutorial “Information Discovery in e-Commerce” taught at SIGIR 2018. We thank the audience for their feedback and questions. We also thank our colleagues Mozhdeh Ariannezhad, Hongshen Chen, Jiawei Chen, Zhumin Chen, Songgaojun Deng, Zhuoye Ding, Yue Feng, Stefan Grafberger, Paul Groth, Yulong Gu, Shuyu Guo, Ziyi Guo, Maria Heuss, Mariya Hendriksen, Na Huang, Sami Jullien, Barrie Kersbergen, Jiahuan Lei, Dongdong Li, Ming Li, Xiang Li, Xinyi Li, Zhenyang Li, Xiaozhong Liu, Hengliang Luo, Si Luo, Yougang Lyu, Jun Ma, Yao Ma, Pengjie Ren, Emma de Rijke, Fatemeh Sarvi, Sebastian Schelter, Xinlei Shi, Clemencia Siro, Hongye Song, Olivier Sprangers, Changlong Sun, Fei Sun, Weiwei Sun, Jiliang Tang, Bart Voorn, Jingang Wang, Shuaiqiang Wang, Xuepeng Wang, Zihan Wang, Long Xia, Xin Xin, Zhen Zhang, Jiashu Zhao, Xiangyu Zhao, and Yihong Zhao for help, feedback, and inspiration. We thank our editors Yiqun Liu and Mark Sanderson for support, patience, and valuable feedback. This research was supported by Alibaba DAMO Academy, Baidu.com, JD.com, and Meituan, as well as AIRLab, a collaboration between Ahold Delhaize and the University of Amsterdam, the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project nr. 024.004.022, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO) and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.